You need to make an aqueous solution of 0.145 m iron(lI) chloride for an experiment in lab, using a 200 grams of water. How much solid iron(II) chloride should you add?y
1
Expert's answer
2019-05-17T03:29:12-0400

Since m(H2O) = 200 g, Vsol ≈ 200 mL = 0.2 L.

n(FeCl2) = C(FeCl2) × V(FeCl2) = 0.145 M × 0.2 L = 0.029 mol.

m(FeCl2) = M(FeCl2) × n(FeCl2) = 127 g/mol × 0.029 mol = 3.683 g.